David Abbot is interested in purchasing a bond Before-tax cost of debt and after-tax cost of debt Personal Finance Problem issued by Sony. He has obtained the following information on the security: Sony Bond Par value $1000 Coupon interest rate 6.5% Cost $930Years to maturity 10 Corporate tax rate 20% Answer the following questions: a. Calculate the before-tax cost of the Sony bond using the bonds yield to maturity (YTM) b. Calculate the after-tax cost of the Sony bond given the corporate tax rate. a. The before tax cost of the Sony bond using the bonds yield to maturity YTM is % Round to two decimal places. b. The after-tax cost of the Sony bond given the corporate tax rate is %. (Round to two decimal places.)
Answer #1

Par Value of the bond = $1000

Coupon Interest Rate = 6.5%

Cost of the Bond = $930

Years to Maturity = 10years

Tax Rate= 20%

Coupon Interest = Par value of the bond * Interest rate = $1000*6.5% = $65

Cost of the bond = Coupon Interest* pvifa (ytm%, 10years) + Par Value of the bond* pvif (ytm%, 10yrs)

$930 = $65* pvifa (ytm%, 10yrs) + $1000* pvif (ytm%, 10yrs)

Now interpolating the results;

Let us suppose YTM% = 9%

$930= $65* 6.418 + $1000* 0.4224= $839.57

Let us suppose YTM% = 7%

$930 = $65* 7.0236 + $1000*0.5083= $964.84

Or, (9-7) %/ (9-r)% = (839.57-964.84)/ (839.57-930)

Or, 2%/ (9-r)% = -125.27/-90.43

Or, 9-r = 1.45%

Or, r = 7.55%

So before tax cost of Sony’s Bond using the YTM (%) is 7.55%

b. After tax Cost of Sony’s Bond using the YTM (%) is Pre-Tax cost of Debt* (1- Tax rate)

Or, Post tax Cost of Sony’s Bond using the YTM (%) is 7.55%*0.8= 6.05%

Pre-Tax Cost of Debt (%) = 7.55%

Post- Tax Cost of Debt (%) = 6.05%